OpenStudy (anonymous):

evaluate sin (-225) and sec 10pi/3

OpenStudy (solomonzelman):

sin(-225)=sin(270-45)=sin(270)cos(45) - cos(270)sin(45) sin(270) =-1 cos(270)= 0 sin(45) and cos(45) are both \(\LARGE\color{blue}{ \bf \frac{2}{\sqrt{2}} }\)

OpenStudy (solomonzelman):

Let me know if you need more help.

OpenStudy (anonymous):

yes I use help with the second one if possible